OpenStudy (anonymous):

Suppose that three computer boards in a production run of forty are defective. A sample of five is to be selected to be checked for defects. How many different samples can be chosen, and how many samples will contain at least one defective board?

OpenStudy (anonymous):

I know how to get the first part, but not sure what method to use for the second

OpenStudy (anonymous):

ie there are 40C5 different samples, right

OpenStudy (anonymous):

am I looking for the total possible combinations of 5 parts with between 1 and 3 parts defective? so, 5c1 +5c2+5c3?

OpenStudy (anonymous):

Is this right?

OpenStudy (anonymous):

40C5 will be the selection this is right..

OpenStudy (anonymous):

Atleast means that there can be one defective or two or three..

OpenStudy (anonymous):

so would it actually be 40c1 +40c2 +40c3?

OpenStudy (anonymous):

So the combinations will be: 1 defective + 4 non defective 2 defective + 3 non defective 3 defective + 2 non defective So combine all this..

OpenStudy (anonymous):

oh, gotcha

OpenStudy (anonymous):

so, 40c1*40c4+40c2*40c3+40c3*40c2?

OpenStudy (anonymous):

So you will get: \[\large ^{3}C_1 \times ^{37}C_4 + ^3C_2 \times ^{37}C_3 + ^3C_3 \times ^{37}C_2 \] In my view...

OpenStudy (anonymous):

because you're selecting from a set of 3 defective ones and 37 good ones

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

need more practice. thank you.

OpenStudy (anonymous):

Do you have answers with you for this??

OpenStudy (anonymous):

I don't, it's an even question in the exercises :( but I think the intuition is right.

OpenStudy (anonymous):

So let me check now if this is right or not..

OpenStudy (anonymous):

there were also some exercises dealing with choosing teams of mixed men and women from a large pool that were solved in a similar fashion, so I think this is right.

OpenStudy (anonymous):

Yeah those and this are based on one concept only..

OpenStudy (anonymous):

So congratulations we are right..

OpenStudy (anonymous):

:) thanks for the help

OpenStudy (anonymous):

Welcome dear...